### v1

# MoDA: Multi-modal Diffusion Architecture for Talking Head Generation

Abstract
Talking head generation with arbitrary identities and speech audio remains a crucial problem in the realm of digital humans and the virtual metaverse. Recently, diffusion models have become a popular generative technique in this field with their strong generation and generalization capabilities. However, several challenges remain for diffusion-based methods: 1) inefficient inference and visual artifacts, which arise from the implicit latent space of Variational Auto-Encoders (VAE), complicating the diffusion process; 2) authentic facial expressions and head movements, resulting from insufficient multi-modal information interaction. In this paper, MoDA handle these challenges by 1) defines a joint parameter space to bridge motion generation and neural rendering, and leverages flow matching to simplify the diffusion learning process; 2) introduces a multi-modal diffusion architecture to model the interaction among noisy motion, audio, and auxiliary conditions, ultimately enhancing overall facial expressiveness. Subsequently, a coarse-to-fine fusion strategy is adopted to progressively integrate different modalities, ensuring effective integration across feature spaces. Experimental results demonstrate that MoDA significantly improves video diversity, realism, and efficiency, making it suitable for real-world applications.
Figure 1: Given a reference image, an audio clip, and other control signals, MoDA excels in producing vivid portrait animation videos in real-time. Beyond basic lip synchronization, it demonstrates strong capabilities in generating diverse facial expressions and natural head movements. Furthermore, it supports fine-grained facial expression control and long-term video synthesis.
1 Introduction
Talking head generation is crucial for the metaverse and a wide range of applications, including immersive telepresence and augmented or virtual reality. This task involves generating a speaking portrait from a single image, guided by the corresponding audio and other auxiliary modalities. Lip synchronization for portrait generation has been extensively studied in computer vision. However, generating realistic and efficient talking head videos remains a persistent challenge.
The early methods [24, 27] based on Generative Adversarial Networks (GANs) [11] were the primary solution for audio-to-portrait translation, demonstrating relatively accurate lip synchronization. However, these methods remain limited in capturing diverse speaking styles, expressive emotions, and natural head movements. Recently, diffusion models [7, 14, 33] have significantly advanced generative modeling, enabling the creation of highly diverse and high-fidelity images and videos. Motivated by this progress, significant research efforts [5, 3, 17, 31, 38, 43, 44, 2, 19] have been made aimed at extending these innovations to the development of portrait image animation.
These methods are typically categorized into single-stage and two-stage frameworks. The former [5, 3, 17, 31, 38, 43] rely on a VAE-style latent space to directly generate video content from the audio input, integrating both motion and appearance in one pass. Despite being conceptually simple and trainable from end to end, two major limitations still persist, as shown in Fig. 2: 1) inefficient inference process. Most diffusion-based approaches must account for not only the diverse mapping of lip movements but also the rendering of visual appearances. This dual consideration significantly increases computational costs, making these methods impractical for real-time applications. 2) Limited capability for long-term video generation and visual artifacts. Methods like EMO [38] and Hallo [43] generate new video content by conditioning on previous frames. These prior frames provide motion information for subsequent generation and simultaneously leak appearance details, leading to the progressive accumulation of pixel errors and deterioration in overall video quality.
The two-stage methods [44, 2, 19] use diffusion models to generate low-dimensional intermediate motion representations, simplifying the learning process. However, the effectiveness of these methods is often compromised by inaccurate representations. This limitation mainly arises from the inherently one-to-many nature of the audio-to-motion task. The commonly used unidirectional cross-attention mechanism fails to support sufficient multi-modal interaction, leading to inconsistent motion sequences when conditioned on arbitrary identities and audio inputs, as discussed in Sec. 4.3.1. Motivated by recent lip-to-speech tasks [40, 26], visual information can provide complementary contextual cues to audio. Sufficient multi-modal interaction helps disambiguate motion generation, resulting in more accurate and identity-consistent outputs.
In this paper, we propose a novel framework named MoDA, designed to generate realistic talking head videos from a single portrait image. MoDA first establishes a joint parameter space that bridges motion generation and neural rendering, consisting of a disentangled motion-appearance representation [13] along with the corresponding audio, emotion, and identity information. As shown in Fig. 3, MoDA introduces Multi-modal Diffusion Transformer (MMDiT) [9], equipped with rectified flow [21], as the backbone diffusion framework to facilitate multi-modal interaction. In this design, MoDA can dynamically adapt audio features based on motion, identity, and emotion, thereby improving the accuracy of motion generation.
However, there are semantic overlaps between modalities within the joint parameter space: for example, audio inherently contains emotional and identity-related cues. Independently learned weights hinder the performance of the model. To facilitate effective feature fusion, a coarse-to-fine fusion architecture is further proposed to progressively enhance feature integration across modalities. Initially, the model maintains modality-specific weights to capture the unique characteristics of each condition. In the intermediate stage, audio and auxiliary conditions, such as identity and emotion, begin to share attention weights. Finally, all modalities are merged into a unified representation space, allowing holistic fusion and improving the overall generative expressiveness. Moreover, to further encourage motion diversity while maintaining temporal consistency, this paper provides an optional Adaptive Lip-motion Synchronization Expert (ALSE), which can be integrated during training.
The contributions of our method can be summarized as follows:
  • This paper proposes MoDA, a novel multi-modal diffusion framework that generates high-fidelity talking head videos from a single portrait image, audio, and additional modalities.
  • A coarse-to-fine fusion strategy is designed to progressively integrate noise motion with audio and other modalities, enabling effective cross-modal interaction.
  • Extensive evaluations on public datasets demonstrate that our method outperforms contemporary alternatives in visual quality and quantitative metrics.
Figure 2: Comparison of our approach with hallo2 [5]. Real-Time Factor (RTF) measures the ratio of inference time to output duration. An RTF <1 indicates that the system can perform real-time generation. Frame-by-frame comparison. Hallo2 shows lower lip synchronization, inference efficiency, and identity consistency than our method.
2 Related Work
2.1 Disentangled Face Representation
In recent years, extensive research has been conducted on disentangled facial representation learning. Some methods utilize sparse keypoints [32, 50] or 3D Morphable Model (3DMM) [1, 20] techniques to explicitly represent facial dynamics. The 3DMM initially projects the 3D head shape into several low-dimensional Principal Component Analysis (PCA) spaces, which provide orthogonal bases. Based on these orthogonal bases, attributes such as identity, pose, and expression can be manipulated in the 3D model using linear blend skinning (LBS). However, these methods may face issues such as inaccurate reconstructions or limited ability to decouple facial attributes. Recent learning-based approaches, such as Face Vid2Vid [42], LivePortrait [13], and Megaportrait [8], have been introduced to disentangle facial representations within a nonlinear parameter space, significantly enhancing the expressive power of these models. Moreover, by leveraging a more compact parameterization, these methods can effectively capture intricate details and dynamic facial expressions, offering greater flexibility in both facial reconstruction and animation.
2.2 Audio-driven Talking Head Generation
In audio-driven digital human technology, one-shot methods have gained considerable attention, enabling the generation of dynamic and expressive avatars from a single image input. Existing methods can generally be classified into two categories: single-stage and two-stage audio-to-video generation. The former [17, 5, 27, 31, 43, 12, 6] directly maps audio features to video frames in an end-to-end manner. In contrast, the latter[19, 44, 2, 35, 51] introduces an intermediate representation, such as motion sequences or keypoints, serving as a bridge between audio and video synthesis.
Early single-stage methods [12, 27, 36] leveraged GANs to focus on lip-sync accuracy while keeping other facial attributes unchanged. Recent advancements have incorporated diffusion-based approaches [43, 31, 5, 38, 3], expanding the scope of research by mapping audio inputs to diverse facial expressions and continuous natural head movements. However, these methods rely on denoising and rendering images within the VAE space, where appearance and motion remain highly entangled, resulting in substantial computational overhead and low inference efficiency.
The two-stage methods address the above limitations by introducing a disentangled facial space as an intermediate representation. Early methods [47, 48] used explicit landmarks or 3DMM as intermediate representations to bridge the gap between audio inputs and animated outputs, allowing for more controllable and interpretable motion synthesis. Recently, VASA-1 [44], Ditto [19], and JoyVASA [2] have shifted towards implicit facial representations [42, 8, 13] while employing DIT-based models for audio-to-motion mapping, leading to more expressive and natural video synthesis. However, relying solely on cross-attention for lip-sync generation neglects the rich multimodal interactions and deep-level information within the input signals, thereby limiting the diversity and expressiveness of the generated outputs.
Figure 3: Overall architecture of the proposed MoDA and illustration of our Motion Generation Network. The appearance extractor $\mathcal{F}$, motion extractor $\mathcal{M}$, warping module $\mathcal{W}$, and decoder $\mathcal{G}$ are frozen. A motion feature generation model based on a Diffusion Transformer is then trained to generate motion features.
3 Method
As illustrated in Fig. 3, MoDA is a two-stage framework. The first stage focuses on disentangling facial attributes, while the second stage generates motion based on audio and other conditions. The final output is rendered using the generated motion features and the 3D appearance features obtained from the first stage. For the first stage, we leverage the LivePortrait [13] framework to achieve fine-grained control over facial attributes, including expression, head pose, and gaze direction. Our primary contribution lies in the second stage, the motion feature generation network, which is built around four key components: Joint Parameter Space (see Sec. 3.2.1), Motion Transformer Blocks (see Sec. 3.2.2), Coarse-to-Fine Feature Fusion (see Sec. 3.2.3), and a carefully designed Loss Function for training (see Sec 3.2.4). These modules work collaboratively to ensure long-term identity consistency and natural facial expression generation.
3.1 Rectified Flow
Denoising Diffusion probabilistic models (DDPMs) [34] have emerged as a powerful framework for generative modeling by formulating the data generation process as an iterative denoising procedure. In the forward diffusion process, Gaussian noise $\epsilon$ is gradually introduced into the data distribution across $T$ discrete timesteps, producing noisy latent features:$z_{t}=\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon,$ where $\alpha_{t}$ represents a variance schedule that determines the noise level at each timestep, and $z_{0}$ is the raw data. The model is trained to reverse this process by taking the noisy latent representation $z_{t}$ as input and estimating the added noise $\epsilon$. The training objective is defined as:$\mathcal{L}=\mathbb{E}_{z_{t},c,\epsilon\sim\mathcal{N}(0,1),t}\left[||\epsilon-\epsilon_{\theta}(z_{t},t,c)||_{2}^{2}\right],$ where $\epsilon_{\theta}$ denotes the noise prediction generated by the model, and $c$ represents additional conditioning signals, such as audio or motion frames, which are particularly relevant in the generation of talking videos.
Recently, Stable Diffusion 3 (SD3) [9] has advanced the paradigm by introducing Rectified Flow [25, 21] that optimizes the traditional DDPMs objective:
$\mathcal{L}=\mathbb{E}_{z_{t},c,\epsilon\sim\mathcal{N}(0,1),t}\left[||(\epsilon-z_{0})-v_{\theta}(z_{t},t,c)||_{2}^{2}\right],$ (1)
where $z_{t}=(1-t)z_{0}+t\epsilon$, and $v_{\theta}$ denotes the velocity field. After the rectified flow training is completed, the transition from $\epsilon$ to $z_{0}$ can be formulated using the numerical integration of an ordinary differential equation (ODE):
$z_{t-\frac{1}{N}}=z_{t}+\frac{1}{N}v_{\theta}(z_{t},t,c),$ (2)
where $N$ denotes the discretization steps of the interval $[0,1]$. The piecewise linear denoising process improves training stability. Furthermore, since our audio-to-motion task does not involve complex pixel information, this approach is particularly well-suited to our needs. Given these advantages, we adopt the rectified flow for training.
3.2 Motion Feature Generation with Diffusion Transformer
Similar to previous approaches [19, 44, 10], the primary conditioning signal for MoDA is the audio feature, incorporating additional auxiliary signals for the speaker’s identity and emotion to enhance controllability. However, instead of simply concatenating these conditioning signals, we propose to enhance the intrinsic identity and emotional information of the audio feature. This approach effectively balances identity and emotion across audio and speakers, enabling more natural lip control in real-world scenarios.
3.2.1 Joint Parameter Space
As shown in Fig. 3, we incorporate the existing facial reenactment framework [13] to extract disentangled facial representations. Specifically, the motion extractor $\mathcal{M}$ yields expression deformations $\delta$, head pose parameters $(R,t)$, the canonical keypoints of the source image $x_{c}$, and a scaling factor $s$. The motion representation $(R_{s},\delta_{s},t_{s},s_{s})\in\mathbb{R}^{70}$, serves as an identity-agnostic representation of the source input and is used to train MoDA to predict $(\hat{R},\hat{\delta},\hat{t},\hat{s})$ given audio input.
$x_{s}$ $=s_{s}\cdot(x_{c}R_{s}+\delta_{s})+t_{s}$ , (3)
$\hat{x}$ $=\hat{s}\cdot(x_{c}\hat{R}+\hat{\delta})+\hat{t}$
Subsequently, the warping field estimator $\mathcal{W}$ computes a field from $x_{s}$ and $\hat{x}$ to deform the 3D features $f_{s}$, which are then passed to the generator $\mathcal{G}$ to synthesize the target image. The audio features $f_{a}$ are extracted using the wav2vec [29] encoder. To maintain consistency in the identity feature space across various scenarios, we use the canonical keypoints of the source image $x_{c}$ as identity information and generate identity features $f_{i}$. For facial emotion signals, we use a visual emotion classifier [28] to extract the speaker’s emotional labels and encode them into corresponding features $f_{e}$. The motion features from the first frame of each clip are used as guiding conditions to ensure inter-frame continuity and generate noisy motion features $f_{m}$. Inspired by EMO2 [37], emotion features $f_{e}$ are added to timestep embeddings to generate timestep features $f_{t}$, which are injected into each motion transformer block by adaptive layer normalization (AdaLN) [23]. AdaLN is used to prevent the degradation of emotion features during the joint-attention operation, ensuring that emotional cues are preserved throughout the interaction.
By integrating these conditioning signals, our model effectively generates realistic and temporally consistent motion features.
Figure 4: The detail of the Motion Transformer Block and the Joint-attention. (a) Cross-attention uses noise as the query and conditions as key and value. (b) Joint-attention projects noise and conditions separately and concatenates them before attention.
3.2.2 Motion Transformer Blocks
As shown in Fig. 4 (c), the Motion Transformer Block consists of three main components: Modality-specific paths, Joint-attention and Rotational Position Encoding.
Modality-specific paths are designed to distinguish representations from different modalities. Specifically, each modality is equipped with its own adaLN and a modulation mechanism [23] to improve the conditional generation capabilities of the model.
Joint-attention is employed to enable interaction across different modalities. All modalities are first projected onto their respective query (Q), key (K), and value (V) representations, which are then concatenated in order along the sequence dimension. The combined sequence is processed through an attention operation, after which the attended features are split back into their respective modalities in the original order.
To enhance temporal alignment between noisy motion and other conditional features, we adopt Rotational Positional Encoding (RoPE) [46] instead of the absolute positional encoding used in MMDiT. Specifically, we first expand the expression and identity features to match the sequence length of the noisy motion and audio features. We then apply aligned RoPE across all modalities, which facilitates better temporal synchronization and more consistent feature representations.
3.2.3 Coarse-to-Fined Feature Fusion
Although assigning modality-specific paths with separate QKV projections and FeedForward Networks (FFNs) in the attention mechanism can enhance cross-modal interaction, this design often overlook the inherent semantic commonalities shared across modalities. Independently learned weights hinder effective multimodal feature fusion and introduce redundant parameters, potentially causing convergence issues. This issue is particularly pronounced in tasks like audio-to-motion generation, where no pixel-level inputs are involved and the semantic gap between modalities is relatively small. In such cases, maintaining entirely separate parameterizations fails to provide meaningful benefits and instead leads to duplicated representations. To enhance the model’s capacity for multi-modal understanding, we introduce a Coarse-to-Fine Feature Fusion and Interaction strategy. This interaction strategy progressively integrates multi-modal features while reducing unnecessary parameters, thereby enhancing the model’s multi-modal understanding and training stability.
As illustrated in Fig. 3, the architecture progresses through three stages: a four-stream, two-stream, and single-stream process. In the four-stream stage, each modality is processed independently to learn modality-specific representations, facilitating early-stage feature differentiation and alignment. In the two-stream stage, the audio stream is concatenated with emotion and identity features to form a merged stream that shares weights. This design encourages a balanced integration of emotional and identity cues from both the audio and the reference image. Finally, all modalities are unified into a single representation stream to allow deeper cross-modal fusion and enhance the generative expressiveness of the model.
Table 1: Comparison with existing methods on the HDTF and CelebV-HQ test sets. ↑ Higher is better. ↓ Lower is better. Best results are in bold, second-best are underlined.
Dataset Method FVD $\downarrow$ FID $\downarrow$ F-SIM $\uparrow$ Sync-C $\uparrow$ Sync-D $\downarrow$ Smooth (%) $\uparrow$
HDTF GT - - 0.860 7.267 7.586 0.9959
EchoMimic 207.987 29.633 0.887 2.744 11.805 0.9939
JoyVASA 229.634 32.584 0.953 5.255 9.600 0.9974
JoyHallo 256.226 44.842 0.852 7.360 7.984 0.9944
Hallo 216.573 34.350 0.878 7.087 7.941 0.9950
Hallo2 229.806 34.426 0.871 7.102 7.976 0.9951
Ours (sync) 191.292 30.449 0.925 8.183 7.065 0.9970
Ours 174.622 28.182 0.927 7.369 7.744 0.9971
CelebV-HQ GT - - 0.861 5.837 7.989 0.9964
EchoMimic 258.451 47.169 0.837 2.610 11.216 0.9946
JoyVASA 271.231 44.574 0.918 5.171 8.632 0.9971
JoyHallo 282.081 57.247 0.813 6.041 8.418 0.9945
Hallo 245.101 44.411 0.851 5.629 8.384 0.9952
Hallo2 242.352 46.615 0.851 5.671 8.397 0.9953
Ours (sync) 205.307 44.201 0.916 6.552 7.635 0.9972
Ours 214.182 42.001 0.913 5.757 8.346 0.9972
Table 2: Comparison with existing methods on the wild test dataset. The best results are bold, and the second are underlined.
Method F-SIM $\uparrow$ Sync-C $\uparrow$ Sync-D $\downarrow$ RTF $\downarrow$
EchoMimic 0.870 2.292 12.130 48.657
JoyVASA 0.924 5.569 9.368 1.717
JoyHallo 0.825 7.000 8.167 59.735
Hallo 0.848 6.051 8.730 59.190
Hallo2 0.849 6.386 8.523 60.305
Ours (sync) 0.896 7.710 7.469 0.846
Ours 0.895 6.862 8.088 0.846
3.2.4 Loss Function
Our loss function utilizes the rectified flow loss and Eq. refeq:2 is rewritten as:
$\mathcal{L}_{RF}=\mathbb{E}_{z_{t},c,\epsilon\sim\mathcal{N}(0,1),t}\left[||x-v_{\theta}(z_{t},t,c)||_{2}^{2}\right],$ (4)
where $x$ represents $\epsilon-z_{0}$, $v_{\theta}$ denotes the velocity field. The velocity loss $L_{vel}$ is introduced to encourage improved temporal consistency:
$\mathcal{L}_{vel}=||x^{\prime}-m^{\prime}||_{2}^{2}+||x^{\prime\prime}-m^{\prime\prime}||_{2}^{2},$ (5)
where $m$ denotes the output of $v_{\theta}(z_{t},t,c)$, and $m^{\prime\prime}$, $m^{\prime\prime}$ denote the first-order and second-order derivatives of $m$. To further enforce lip-sync accuracy, we design the ALSE pretrained in audio and motion features and compute the loss $\mathcal{L}_{ALSE}$. It is worth noting that supervision from the pretrained ALSE is optional and can be applied as needed. In summary, the final loss can be expressed as follows:
$\mathcal{L}$ $=\mathcal{L}_{RF}+\mathcal{L}_{vel}+\lambda_{sync}\cdot\mathcal{L}_{ALSE}$ , (6)
$\lambda_{sync}$ =$=\begin{cases}1,&\text{if}\ \mathcal{L}_{ALSE}<\tau\\ 0,&\text{otherwise}\end{cases}$ , (7)
where $\tau$ is set to 0.4 and controls when the synchronization loss is activated, ensuring that lip-motion supervision is applied only after basic motion diversity has been established. Further details about the Adaptive Lip-motion Synchronization Expert (ALSE) can be found in the supplementary materials.
3.2.5 Classifier-free guidance (CFG)
In the training stage, we randomly assign each of the input conditions, and during inference, we perform the following:
$\hat{z}^{0}=(1+\sum_{c\in C}\lambda_{c})\cdot v_{\theta}(z^{t},t,C)-\sum_{c\in C}\lambda_{c}\cdot v_{\theta}(z^{t},t,C|c=\emptyset),$ (8)
where $\lambda_{c}$ is the CFG scale of condition c. $C|c=\emptyset$ denotes that the condition c is $\emptyset$.
3.2.6 Realtime Inference
The real-time conversational scenario is enabled by low-latency motion generation. We align audio features with the video frame rate and segment the audio into continuous 100-frame chunks for streaming generation. Additionally, leveraging low-dimensional intermediate representations and the rectified flow sampling approach, we reduce DiT inference denoising steps from 40 to 10 while maintaining comparable quality. For detailed comparison, please refer to the supplementary material.
4 Experiments
4.1 Experiment Settings
4.1.1 Dataset and Metrics
Our motion generation model is primarily trained on three publicly available datasets: HDTF [52], CelebV-Text [49], and MEAD [41]. For evaluation, we conducted experiments on three distinct test sets. The first two are derived from public datasets, CelebV-HQ [53] and HDTF, each consisting of 50 randomly sampled clips ranging from 3 to 10 seconds in length. The third is an in-the-wild set with 20 diverse cases, including real individuals, animated characters, dynamic scenes, and complex headwear. Each sample is paired with audio that is speech, emotional dialogue, or singing. We utilize several evaluation metrics to assess the performance of the proposed method. The Fréchet Inception Distance (FID) [30] and the Fréchet Video Distance (FVD) [39] are used to assess the quality of the generated output, while the F-SIM [38] measures facial similarity. In addition, sync-C [4] and sync-D [4] metrics are introduced to evaluate lip synchronization between different methods. A temporal smoothness metric [16] is also utilized to monitor the continuity of generated motion.
4.1.2 Implementation Details
During training, we randomly sample 80-frame segments from video clips to train the motion generation model. The model is trained for approximately 500 epochs on 8 NVIDIA H20 GPUs with a batch size of 512, using the Adam optimizer with a learning rate of 1e-4. During training, we apply a dropout probability of 0.1 for each emotion condition, while the dropout probability for speech is set to 0.5. Furthermore, the model is structured with 3 four-stream blocks, 6 two-stream blocks, and 12 single-stream blocks to progressively enhance multi-modal feature integration.
4.2 Results and Analysis
We juxtapose the quantitative results of the proposed method against those of EchoMimic [3], JoyVASA [2], JoyHallo [31], Hallo [43], and Hallo2 [5]. We also compare two variants of our method: Ours (MoDA without ALSE) and Ours (Sync) (MoDA with ALSE).
Figure 5: Qualitative Comparisons with State-of-the-Art Method. As single-frame images cannot fully represent synchronization, naturalness, and stability, we provide complete video comparisons in the supplementary materials.
Figure 6: Generation results for portraits and audio in diverse styles are presented. We also demonstrate long-video inference and fine-grained control over facial expressions.
4.2.1 Quantitative Comparison
The quantitative results on the CelebV-HQ test and HDTF dataset are shown in Table 1. On the HDTF dataset, MoDA consistently outperforms all existing methods across all evaluation metrics. In particular, our method achieves the lowest FID and FVD scores, outperforming the second-best methods by 4.9 % and 16.0 %, respectively. This demonstrates the superiority of our method in terms of visual naturalness and the overall quality of the generated frames. The disentanglement of appearance and motion enables our method to better preserve identity consistency compared to existing approaches, as evidenced by the second-highest F-SIM score. Notably, JoyVASA achieves the highest F-SIM, likely due to its relatively constrained head movements, which contribute to higher frame-wise structural similarity.
On the CelebV-HQ test dataset, our method consistently outperforms the five baseline methods in all metrics except Sync-D, highlighting its robustness and effectiveness. Although MoDA slightly underperforms JoyHallo in synchronization confidence (Sync-C), it achieves a notable improvement in synchronization distance (Sync-D), along with further gains in motion smoothness. A detailed analysis can be found in Sec. 4.2.2. Table 1 also reports the results with and without incorporating ALSE. Introducing ALSE to supervise the synchronization between audio and keypoints leads to a significant improvement in lip-sync metrics.
Table 2 compares the performance of our method with existing approaches on the wild test dataset, which features more diverse characters and complex backgrounds. Our method consistently outperforms others on multiple metrics, demonstrating its robustness in unconfined scenarios. In addition, we evaluate the inference performance using the real-time factor (RTF). RTF < 1 indicates that the system meets the basic requirement for real-time inference. Our method achieves competitive efficiency, making it suitable for real-world applications that demand both quality and speed.
4.2.2 Qualitative Evaluation
As shown in Fig. 5, we selected two types of cases from the wild dataset for visual comparison. For each character, we generated videos using each method and selected frames from the same location for comparison. Analyzing the results, previous one-stage-based methods, including Hallo2, Hallo, and JoyHallo, suffer from appearance blurring and unnatural expressions during temporal inference due to the strong entanglement between appearance and motion. In contrast, our proposed two-stage framework effectively mitigates these issues, ensuring high consistency in the generated details. Compared to the two-stage method JoyVASA, MoDA generates richer expressions, better lip synchronization, and more natural head movements, thanks to our multi-modal motion generation network, which effectively integrates deep information across different modalities.
Table 3: Ablation study results on major architectural choices.
Method FVD $\downarrow$ Sync-C $\uparrow$ Sync-D $\downarrow$
w/o C2F 221.631 5.535 8.465
w CABA 232.291 5.331 8.692
Full Model 214.182 5.757 8.346
4.2.3 Visualization Results in Complex Scenarios.
We further investigate MoDA’s generation performance in complex scenarios. Specifically, for the visual modality, we utilize portraits of both real humans and animated characters, each paired randomly with various audio types, including speech, singing, recitation, and others. As shown in Fig. 6, our method demonstrates strong performance in various complex scenarios. In addition, it is capable of efficient long-duration inference and fine-grained facial expression control.
4.3 Ablation Study
To validate the effectiveness of our method components, we deactivate each of them and report result in Table 3. For more ablation experiments and implementation details, please refer to the supplementary materials.
4.3.1 Cross-attention
We replace our model with a Cross-Attention-Based Architecture (CABA) to perform multi-modal fusion for evaluation purposes. As shown in Table 3, this substitution results in a performance drop in all metrics. Moreover, as illustrated in Fig. 7, the subject exhibits mouth-closing failures when lacking deep cross-modal interaction. To investigate this, we conducted two experiments under the same settings used in the w/ CABA variant as shown Fig. 7: 1) replacing the audio with the same image did not resolve the mouth closure issue (Replace audio); 2) replacing the image with the same audio addressed the issue (Replace image). These results suggest that cross-attention fails to adapt audio features to different identity conditions. In contrast, MoDA enables dynamic adaptation of audio features based on contextual factors such as motion, identity, and emotion.
4.3.2 Coarse-to-Fined Feature Fusion and Interaction
We replace the Coarse-to-Fine Feature Fusion with a fully four-stream architecture (w/o C2F). As shown in Table 3, this design leads to performance degradation, particularly in synchronization and motion quality, emphasizing the necessity of progressive feature fusion for effective cross-modal integration. These results indicate that independently learning weights introduces substantial redundant parameters, resulting in convergence challenges (with 904M parameters in the w/o C2F variant). In contrast, by gradually sharing weights, the proposed coarse-to-fine strategy significantly reduces the parameter count to 370M, leading to both greater efficiency and improved performance.
4.4 Discussion
A key limitation of our framework lies in the performance constraints introduced by the first-stage model, including the appearance encoder, motion encoder, and decoder. In particular, the Liveportrait model used at this stage performs poorly when handling large pose variations, leading to a degradation in output quality. Moreover, it often produces unnatural distortions when processing subjects with complex head accessories. These issues may arise from limitations in the training data or intrinsic constraints of the GAN-based generator architecture.
To address these issues, future work will explore more robust first-stage generation techniques. We plan to improve the training dataset by incorporating a wider range of head pose variations and complex head accessories. In addition, we will investigate more advanced video generation models [45, 18, 15, 22] to further improve the quality and expressiveness of the result.
Figure 7: Visualization of ablation studies on key components.
5 Conclusion
We present MoDA, a two-stage multimodal diffusion architecture specifically designed for one-shot talking-head synthesis. Our method effectively leverages multimodal information to map audio signals to motion sequences in an identity-agnostic latent space, followed by a dedicated face generator to produce high-quality videos. MoDA can synthesize natural and coherent talking head videos from a single portrait image with high efficiency, while also supporting fine-grained expression control. It significantly outperforms previous state-of-the-art methods in terms of video quality, motion diversity, and naturalness.